2. Find the Fourier transform of signal z(t) shown below. z(ty
Answer #1

signal 3ct) is that the Given a - St- 3 (t) = 18+) 2 t - 18 ه د (+)a 3(t) = 5 - 25 t< ا ا 2 | 2- - Jourier traustorm alt) the